OpenStudy (anonymous):

a 1= 18, d = –10, n = 12

OpenStudy (amistre64):

looks like some information about sequences .... arithmetic maybe?

OpenStudy (anonymous):

it is A.P

OpenStudy (amistre64):

a1 = 18 a2 = 18 - 10 a3 = 18 - 10 - 10 a3 = 18 - 10 - 10 - 10 ... a12 = 18 - .........

OpenStudy (anonymous):

no just regular, i just need an answer

OpenStudy (amistre64):

we are not here to "just give an answer" we are here to help you learn and study concepts

OpenStudy (anonymous):

whats a formula i can use for these problems

OpenStudy (amistre64):

brute math is one approach. just keep subtracting 10 from to get to the next term ... until you get to the a12 term

OpenStudy (anonymous):

an=a+(n-1)d

OpenStudy (anonymous):

simply solve from this

OpenStudy (amistre64):

we can also develop a formula a1 = 18 a2 = 18 - 10 = a1 - 10 a3 = 18 - 10 - 10 = a2 - 10 = a1 - 10 - 10 = a1 -10(2) a4 = 18 - 10 - 10 - 10 = a3 - 10 = a1 -10(2) - 10 = a1 - 10(3) ... the pattern that develops is: a1 = 18 - 10(0) a2 = 18 - 10(1) a3 = 18 - 10(2) a4 = 18 - 10(3) ... an = 18 - 10(n-1)

OpenStudy (anonymous):

how do i know hwat "a' is

OpenStudy (amistre64):

"a" is not a variable, its a name ...

OpenStudy (amistre64):

a1 defines the first term, or element, of the sequence of A A = {a1, a2, a3, a4, a5, ..., an, ...} the rule for A is: an = a1 -10(n-1)

OpenStudy (anonymous):

so with the formula an=a+(n-1)d what do i put for a

OpenStudy (amistre64):

that was the other persons shorthand for stating the first term: a1= 18, d = –10, n = 12 ^^^^^^ they used "a" instead of "a1"

OpenStudy (anonymous):

im still jsut not getting the right answer

OpenStudy (amistre64):

im not sure what answer you are getting .... can you post how you are workng it?

OpenStudy (anonymous):

an=a+(n-1)d an+18+(12-1)-10 an+18-110

OpenStudy (amistre64):

thats a fair assessment but lets clean that up some, since n=12 \[\large a_{12}=18-10(12-1)\] \[\large a_{12}=18-110\] what is 18 - 110?

OpenStudy (anonymous):

-92

OpenStudy (amistre64):

then without any other information to go by ... -92 would be what you are trying to obtain

OpenStudy (anonymous):

the answer choices are -552,-888,768, and -444

OpenStudy (amistre64):

yeah, thats odd. can you post a screenshot?

OpenStudy (anonymous):

idk how to do that on a computer , but doesnt that make no sense?

OpenStudy (amistre64):

there are a few keys that selodom get used: prnt scrn, scroll lock, pause .... the prnt scrn button copies whatever is on your monitor paste that into a paint file and save; then attach the file with the attach file button at the bottom of this screen

OpenStudy (amistre64):

or paste it into a word document ... and save

OpenStudy (anonymous):